OpenStudy (anonymous):

in lecture 16 on the elastic and inelastic collisions, the prof asks how can a wall have a momentum and no kinetic energy... can anybody help me understand that please?

OpenStudy (anonymous):

I think your professor mean that wall have a momentum although small, so have forces from f=dp/dt, and then why no kinetic energy, i think because Ek=1/2mv^2 than p=mv, because of square so Kinetic energy negligible, energy transferred to other energy like sound